Solve the following system by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
x+5y=10
-2x-10y=-20

Answer by vertciel(183)

x + 5y = 10 --(1)
-2x -10y = -20 --(2)
(1) x 2 = 2x + 10y = 20 --(3)
(3) + (2) = 0
The solution to these two equations is 0. I'll let you figure out why.